# Mohamed Bouadi, Pratinav Seth Aditya Tanna, Vinay Kumar Sankarapu Lexsi Labs, France & India

### v1

@@ 1  Introduction
Tabular data remain the most common form of real-world data, spanning domains such as healthcare, finance, and scientific research. Despite the remarkable progress of deep learning in natural language processing [liu2023pre, yu2024natural] and computer vision [goldblum2023battle], gradient boosted trees (GBTs) remain the predominant state-of-the-art (SOTA) for tabular prediction tasks. In other data modalities, foundation models—particularly Large Language Models (LLMs) [DBLP:conf/nips/00020D24, DBLP:conf/nips/LiZLML0HY24]—have significantly advanced the ability to tackle new tasks and few-shot learning. This is largely due to their remarkable in-context learning (ICL) capabilities [DBLP:conf/iclr/Zhang0YOKZY025, DBLP:conf/uss/CarliniTWJHLRBS21], which enable them to capture patterns directly from prompts without updating their parameters. This success combined with the pervasiveness of tables have spurred interest in tabular foundation models [DBLP:conf/icml/BreugelS24].
@@
Although LLMs are primarily designed to process natural language, recent efforts have explored fine-tuning them for tabular data tasks [tabllm, tmlr]. These approaches typically rely on table serialization, which is the process of converting table rows into text or sentences suitable for tokenization. For instance, [DBLP:conf/nips/0001PS24] fine-tuned a Llama 3-8B model on a large corpus of serialized tables and demonstrated that this strategy can outperform traditional tree-based models in few-shot scenarios. However, such language model–based approaches face inherent challenges. Their limited context windows restrict the number of serialized examples that can be processed simultaneously (e.g., up to 32 or 64 shots in [DBLP:conf/nips/0001PS24]), and it remains uncertain whether LLMs can reliably interpret and reason over numerical values [DBLP:conf/naacl/ThawaniPIS21].
@@
Recently, tabular in-context learning has emerged, adapting the ICL paradigm—central to large language models—to tabular data, enabling pretraining across diverse tables and rapid task adaptation without gradient updates. TabPFN [tabpfn] pioneered this approach by meta-training a transformer on synthetic datasets generated via structural causal models. Its encoder–decoder design lets test samples attend to training examples for zero-shot prediction, but alternating column- and row-wise attentions make large training sets computationally costly. TabDPT [tabdpt] achieves comparable performance using similarity-based retrieval, though its diffusion process adds overhead. TabPFN-v2 [tabpfn2] extended row-based encoding to datasets exceeding 10,00010{,}000 samples. TabICL [tabicl] further refined this with a table-native transformer comprising column embeddings, row interactions, and an ICL head. Its SetTransformer-based column encoder and label-aware in-context learner achieve state-of-the-art results, but its row encoder (tf_row) applies a single homogeneous attention over all features, ignoring local groupings, multi-scale interactions, and structured aggregation. This limitation is pronounced in high-dimensional, heterogeneous tables where features naturally form semantic groups (e.g., demographics, vitals, labs) and dependencies span multiple scales. Additionally, TabICL treats synthetic tables as individual supervised tasks rather than support/query episodes, producing an implicit few-shot signal that may misalign with test-time objectives.
@@ 2  Orion-Bix: Proposed Approach
Orion-Bix retains two core TabICL [tabicl] components: the SetTransformer column encoder and the label-aware in-context learner, while introducing two major changes: (1) a biaxial row encoder that replaces the single-stage tf_row and provides structured multi-scale feature reasoning, and (2) a meta-learning training regime that constructs explicit support/query episodes from synthetic tables, optionally using kNN-based support selection. An overview of the complete architecture is shown in Figure 1.
@@ 2.2  Column-wise Embedding
We adopt the original TabICL [tabicl] column-wise embedder to map each feature to a $d$-dimensional representation per cell. The embedder uses a Set Transformer, $TF_{\text{col}}$, that treats each column as a permutation-invariant set of values across all rows.
@@ 2.5  Meta-Training with Synthetic Episodic Data
In summary, Orion-Bix keeps the core strengths of TabICL [tabicl]—column-wise SetTransformer embeddings and a label-aware in-context Transformer—but introduces a biaxial row encoder that better matches the structure of tabular feature spaces and a meta-learning training regime that more directly optimizes for few-shot in-context performance.
@@ 3  Experimental Evaluation
We evaluate Orion-Bix against TabICL and other baselines to highlight the benefits of biaxial attention and meta-learning. Our experiments focus on three aspects: (i) domain-specific performance on datasets with natural feature structure, (ii) support set quality to assess robustness, and (iii) few-shot learning curves across varying support sizes. These analyses target the improvements motivating Orion-Bix: structured multi-scale feature interactions and enhanced few-shot adaptation.
@@ 3.1  Evaluation Setup
We construct domain-specific evaluation suites by grouping datasets from public benchmarks (e.g., TALENT [talent] and OpenML-CC18 [openmlcc18]) according to application domain, allowing assessment in contexts where feature structure and heterogeneity are most relevant. The domains are: Medical Finance. These domains exhibit the high-dimensional, structured, and multi-scale features that Orion-Bix is designed to handle.
@@
For each dataset, we report overall classification accuracy (ACC), class-weighted F1 to account for imbalance, and mean rank across datasets within each domain based on accuracy. All models use official train/test splits unless noted.
@@ 4  Conclusion
In summary, Orion-Bix preserves the core strengths of TabICL while adding a biaxial row encoder and an episodic meta-learning schedule that directly targets few-shot in-context performance.

### v2

@@ 1  Introduction
Tabular data remain the most common form of real-world data, spanning domains such as healthcare, finance, and scientific research. Despite the remarkable progress of deep learning in natural language processing [11, 18] and computer vision [6], gradient boosted trees (GBTs) remain the predominant state-of-the-art (SOTA) for tabular prediction tasks. In other data modalities, foundation models—particularly Large Language Models (LLMs) [20, 10]—have significantly advanced the ability to tackle new tasks and few-shot learning. This is largely due to their remarkable in-context learning (ICL) capabilities [19, 3], which enable them to capture patterns directly from prompts without updating their parameters. This success combined with the pervasiveness of tables have spurred interest in tabular foundation models [16].
@@
Although LLMs are primarily designed to process natural language, recent efforts have explored fine-tuning them for tabular data tasks [7, 4]. These approaches typically rely on table serialization, which is the process of converting table rows into text or sentences suitable for tokenization. For instance, [5] fine-tuned a Llama 3-8B model on a large corpus of serialized tables and demonstrated that this strategy can outperform traditional tree-based models in few-shot scenarios. However, such language model–based approaches face inherent challenges. Their limited context windows restrict the number of serialized examples that can be processed simultaneously (e.g., up to 32 or 64 shots in [5]), and it remains uncertain whether LLMs can reliably interpret and reason over numerical values [15].
@@
Recently, tabular in-context learning has emerged, adapting the ICL paradigm—central to large language models—to tabular data, enabling pretraining across diverse tables and rapid task adaptation without gradient updates. TabPFN [8] pioneered this approach by meta-training a transformer on synthetic datasets generated via structural causal models. Its encoder–decoder design lets test samples attend to training examples for few-shot prediction, but alternating column- and row-wise attentions make large training sets computationally costly. TabDPT [12] achieves comparable performance using similarity-based retrieval, though its diffusion process adds overhead. TabPFN-v2 [9] extended row-based encoding to datasets exceeding 10,00010{,}000 samples. In addition, OrionMSP[2] introduced new architecture that is explicitly designed around feature/sample structure and mixed datatypes using block sparse attention. TabICL [13] proposed a table-native transformer comprising column embeddings, row interactions, and an ICL head. Its SetTransformer-based column encoder and label-aware in-context learner achieve state-of-the-art results, but its row encoder (tf_row) applies a single homogeneous attention over all features, ignoring local groupings, multi-scale interactions, and structured aggregation. This limitation is pronounced in high-dimensional, heterogeneous tables where features naturally form semantic groups (e.g., demographics, vitals, labs) and dependencies span multiple scales. Additionally, TabICL treats synthetic tables as individual supervised tasks rather than support/query episodes, producing an implicit few-shot signal that may misalign with test-time objectives.
@@ 2  Orion-Bix: Proposed Approach
Orion-Bix retains two core TabICL [13] components: the SetTransformer column encoder and the label-aware in-context learner, while introducing two major changes: (1) a biaxial row encoder that replaces the single-stage tf_row and provides structured multi-scale feature reasoning, and (2) a meta-learning training regime that constructs explicit support/query episodes from synthetic tables, optionally using kNN-based support selection. An overview of the complete architecture is shown in Figure 1.
@@ 2.2  Column-wise Embedding
We adopt the original TabICL [13] column-wise embedder to map each feature to a $d$-dimensional representation per cell. The embedder uses a Set Transformer, $TF_{\text{col}}$, that treats each column as a permutation-invariant set of values across all rows.
@@ 2.5  Meta-Training with Synthetic Episodic Data
In summary, Orion-Bix keeps the core strengths of TabICL [13]—column-wise SetTransformer embeddings and a label-aware in-context Transformer—but introduces a biaxial row encoder that better matches the structure of tabular feature spaces and a meta-learning training regime that more directly optimizes for few-shot in-context performance.
@@ 3  Experimental Evaluation
We evaluate Orion-Bix against TabICL and other baselines to highlight the benefits of biaxial attention and meta-learning. We run all methods via TabTune [14], which standardizes preprocessing and training/evaluation protocols for reproducibility. Our experiments focus on three aspects: (i) domain-specific performance on datasets with natural feature structure, (ii) support set quality to assess robustness, and (iii) few-shot learning curves across varying support sizes. These analyses target the improvements motivating Orion-Bix: structured multi-scale feature interactions and enhanced few-shot adaptation.
@@ 3.1  Evaluation Setup
We construct domain-specific evaluation suites by grouping datasets from public benchmarks (e.g., TALENT [17] and OpenML-CC18 [1]) according to application domain, allowing assessment in contexts where feature structure and heterogeneity are most relevant. The domains are: Medical Finance. These domains exhibit the high-dimensional, structured, and multi-scale features that Orion-Bix is designed to handle.
@@
For each dataset, we report overall classification accuracy (ACC), class-weighted F1 to account for imbalance, and mean rank across datasets within each domain based on accuracy. All models use official train/test splits unless noted. In-context learning models (TabICL, Orion-Bix, TabPFN, TabDPT, Mitra, ContextTab) follow the standard protocol: support examples are provided, and predictions on the test set are made without gradient updates.
@@ 4  Conclusion
Orion-Bix advances tabular in-context learning by strengthening how a model represents row-level structure and how it is trained for few-shot use. Specifically, it replaces TabICL’s flat row encoder with a biaxial row encoder that explicitly mixes (i) full cross-feature attention, (ii) grouped local attention, (iii) hierarchical cross-partition attention, and (iv) relational global attention, and then aggregates these complementary interactions through multi-CLS summaries to preserve multiple aspects of each row representation. This design directly targets common failure modes of tabular transformers in high-dimensional tables, where features naturally form groups and dependencies span multiple scales, while avoiding the information bottleneck of collapsing everything into a single pooled vector.
In addition, Orion-Bix reframes training around explicit episodic meta-learning. Rather than treating each synthetic dataset as one supervised task, training constructs many support/query episodes from synthetic tables via an episode generator (optionally with kNN-based support selection), aligning optimization with the exact inference setting used at test time. This episode-based regime is designed to improve data efficiency and robustness in the few-shot regime by exposing the model to a broader variety of support/query configurations and by optimizing directly for in-context generalization.
Empirically, Orion-Bix shows consistent gains in domains where structured feature interactions matter. On domain-grouped benchmarks, it improves mean rank over TabICL in domains such as Medical and Finance while remaining competitive with strong tree-based baselines. Its advantages are most pronounced in the low-shot setting
4.1  Limitations
Despite these improvements, several limitations remain:
  1. 1.
    Evaluation scope is primarily classification-centric: The experiments emphasize classification metrics such as accuracy and class-weighted F1. Further validation on regression, ranking, calibration, and decision-focused metrics would better characterize real deployment performance.
  2. 2.
    Support-selection benefits are limited: Although the training procedure allows kNN-based episode construction, inference-time results indicate that uniform sampling can slightly outperform kNN-based selection, suggesting curated support selection is not a guaranteed win and may be sensitive to dataset geometry or distance metrics.
  3. 3.
    Potential computational overhead and tuning complexity: The biaxial encoder uses multiple attention passes (standard, grouped, hierarchical, relational) and multi-CLS aggregation. While this improves representational structure, it may increase implementation complexity and require careful tuning (e.g., number of groups, number of CLS tokens, episode size) to balance accuracy and efficiency across datasets and hardware.
  4. 4.
    Large label spaces rely on hierarchical routing: Orion-Bix handles tasks where the number of classes exceeds the native decoder capacity via hierarchical label grouping and tree-based prediction. This introduces another layer where suboptimal grouping could affect accuracy or calibration, especially for fine-grained or highly imbalanced label sets.